OpenStudy (anonymous):

Atrap and Bracken are two rival insurance companies. Atrap and Bracken have premiums of $150 and $100 and deductibles of $2,500 and $3,500 respectively. The average expense of surgery is $25,000. If after 5 years, 25 of the 1,000 people registered with Bracken have gone through surgery, where does Bracken stand in terms of gains or losses?

OpenStudy (anonymous):

choices are $41,100, $39,700, −$36,600, −$37,500, or −$43,000

OpenStudy (anonymous):

5 year income = 5(100)(1,000) = $500,000 Expense = 25(25,000 - 3500) = $537,500 Bracken has lost $37,500 over 5 years.

OpenStudy (anonymous):

You are playing a dice game called Noluz against one other person. Each time you play, both players pay a quarter. You win the money if a 1 or 4 shows face up on the die. The other player wins if a 2 or 5 comes up, and the piggy bank gets the money if a 3 or 6 comes up. This money will then be donated to a charity that you pick.

OpenStudy (anonymous):

What is the expected value per turn for playing Noluz?

OpenStudy (anonymous):

Options??

OpenStudy (anonymous):

Sorry! Options are $0.50, −$0.17, −$0.25, −$0.08, or $0.06

OpenStudy (anonymous):

$0.50 I think, Because each player is giving a quarter and a quarter= 25 cents. 2 people are playing so 25+25=.50 and that is per game.

OpenStudy (anonymous):

Oh okay. thank you so much!!

OpenStudy (anonymous):

Your welcome! Just tag me if you ever need help!

OpenStudy (anonymous):

Okay thanks!